# Principles for Optimizing Quantum Transduction in Piezo-Optomechanical Systems

### v2

@@ Abstract
Two-way microwave-optical quantum transduction is essential to connecting distant superconducting qubits via optical fiber, and to enable quantum networking at a large scale. In Blésin, Tian, Bhave, and Kippenberg’s article, “Quantum coherent microwave-optical transduction using high overtone bulk acoustic resonances” (Phys. Rev. A, 104, 052601 (2021)), they lay out a quantum transducer from microwave photons into telecom-band photons by way of converting to and from GHz-band phonons with piezoelectric, and optomechanical interactions, respectively. In this work, we discuss these coupling interactions from first principles in order to discover what device parameters matter most in determining and optimizing the efficiency of this kind of transducer in the presence of noise.
In addition, we address the post-transduction challenge of separating single photons of transduced light from the classically bright optical pump driving the transducer only a few GHz away in frequency. We first confront this challenge directly, discussing narrow linewidth filtering in integrated optics, and sources of confounding light in the transduction band. That said, we discover that this challenge can be entirely circumvented by utilizing a novel phonon-photon four-wave-mixing optomechanical coupling mechanism that yields transduced light separated from the pump by nearly a full octave in frequency by consuming pairs instead of single pump photons.
@@ I Introduction
Quantum transduction (i.e., where quantum information is transferred from one medium to another) is both a critical and fundamental process needed to connect different species of qubits in a quantum network. This makes more diverse heterogeneous quantum networks possible, but also allows us to leverage the complementary advantages [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] of different species of qubits. In particular, superconducting qubits, which function at microwave frequencies, enable rapid and deterministic gate operations, while photon-based qubits allow one to send quantum information long distances over optical fiber at ambient temperatures with relatively low loss (e.g., $0.18$dB/km for SMF28-Ultra 200 optical fiber at $1550$nm [11]) compared to sending microwave signals over coaxial cables (of the order $10^{2}$ to $10^{3}$dB/km). However, the (microwave/GHz scale) frequencies of light that interact best with superconducting qubits are five orders of magnitude below the (telecom) frequencies at which light propagates best in fiber optics. Moreover, the amount of random thermally generated microwave photons that happen to be in a propagating microwave mode compared to a single-photon microwave signal is large enough that resolving a transduced microwave signal from noise requires a cryogenic operating environment  [^1] As a figure of merit, the expected number of $3.285$ GHz microwave photons in a single mode at $298$K, assuming Boltzmann statistics, would be about $1.89\times 10^{3}$, compared to about $0.26$ at $100$mK.. Unlike classical networking, quantum signals cannot be simply amplified for long-distance transmission due to the no-cloning theorem [13]. Linear amplification of quantum signals (possibly written on individual photons) cannot be carried out without adding noise [14, 15]. This noisy amplification to compensate for loss eventually reaches an irreparable threshold where specific and distinct quantum states (e.g., non-Gaussian states) cannot be reliably transmitted over long distances. In order to connect distant superconducting qubits — as is needed to establish a network of superconducting quantum computers — it is necessary to coherently convert microwave-photons to telecom photons and back again with as little loss as possible.
@@
In this work, we discuss the essential elements underlying the piezoelectric and optomechanical interactions in order to determine what factors are most easily adjusted to maximize the efficiency of transduction. First, we give a brief overview of the system in question, showing how the transduction efficiency is obtained as a function of the coupling interactions. Next, we review the piezoelectric interaction in detail to derive its quantum interaction hamiltonian similarly to what was accomplished in [16]. Following this, we present a review of the optomechanical interaction, deriving the optomechanical coupling constant in terms of an effective photoelasticity tensor, as well as discussing the basic factors underlying the optical waveguide coupling constant. With the fundamentals covered, we explore how changing these factors affects both the peak transduction efficiency as well as its bandwidth, and determine how one may optimize the efficiency through both design and the selection of favorable materials. We conclude by discussing practical challenges to obtaining near-unity transduction efficiency, and strategies toward their resolution. In particular, we address the challenge of cleanly separating the transduced light at single-photon levels of power separated only a few GHz in frequency away from mW-scale-power driving laser light. As will be discussed, we propose a novel optomechanical coupling mechanism using photon-phonon four-wave mixing so that the transduced light is at approximately half the wavelength of the driving pump light instead of very nearly equal to it (i.e., instead of differing by approximately one part in 50,000).
II Background: The Concept of quantum transduction
In order to understand how quantum transduction facilitates quantum networking between different species of qubits, it helps to consider the following:
Ideally, quantum transduction does not change the quantum state of a system per se; instead it changes the identities of the parties comprising the system.
Let us consider storing an arbitrary entangled $N$-qubit state on a collection of superconducting-circuit qubits (which we will abbreviate either as superconducting qubits or as SC qubits) on a single SC-based quantum computer, and that we want to distribute this state so that it is shared among $N$ distant SC-based quantum computers. In order to distribute this entanglement we must do the following (See Fig. 1 for diagram).
  • First, we must couple each of these qubits to microwave transmission lines so that through this interaction, the $N$-party state is now of $N$ distinct propagating microwave photons.
  • Then, using separate transduction platforms for each microwave photon, we can convert the $N$-party state to be one composed of $N$ propagating telecom photons in different fiber optics.
  • Next, we use transduction platforms at the distant nodes to convert the $N$-party state to be one of $N$ microwave photons propagating in transmission lines at their respective nodes
  • Finally, in each of these final nodes, we couple these transmission lines to respective SC qubits, so that we end up with an $N$-party state, distributed among $N$ distantly separated qubits.
Figure 1: Diagram of entanglement distribution via quantum transduction. From left to right four entangled SC qubits (entanglement represented by blue cloud) in a quantum computing node (black rectangle) are transformed into microwave photon qubits propagating in microwave transmission lines, then transduced into telecom-band photons propagating into optical fiber toward distant quantum computers, at which point, the process is reversed to remotely entangle distant SC qubits on separate quantum computers. In principle, one of the SC qubits in the first quantum computer can be stored, so that only the remaining three are distributed to distant SC quantum nodes.
Using only the first half of this protocol, we may use transduction as a source of generating $N$-partite entangled photon states on demand, as prepared on $N$ SC qubits. To date, developing a transducer between photons of microwave and optical (telecom) frequencies of high enough efficiency and low enough noise to make protocols such as these feasible has remained a persistent challenge.
As a fundamental interaction mixing three or more modes, quantum transduction has been realized electromechanically through the Pockels and photoelastic effects; been realized optomechanically through coupling optical and microwave cavities with a movable mirror; and also piezo-optomechanically using the mechanisms described in this paper. We focus on the piezo-optomechanical transducer in [16] because it shows the greatest promise to achieve near-unity transduction efficiency over a broad bandwidth of input photons.
In what follows throughout this paper, we describe what is physically necessary to maximize this microwave-optical transduction efficiency in piezo-optomechanical quantum transducers, with a particular focus on the system in Refs. [16, 17]. That said, the microwave-optical transduction method described here may be used for more than an interface exclusively between distant superconducting qubits, but also connecting SC qubits with a variety of other qubit species (e.g., trapped-ion, nitrogen vacancy centers, neutral atoms, etc) and therefore those disparate species to each other.
III The Piezo-optomechanical transducer
III.1 The physical system
@@
In the piezoelectric stage of the transduction process, there are important elements of design to consider for the microwave frequencies under consideration. We will want to maximize the amount of microwave (electromagnetic) energy being converted into mechanical (vibrational) energy, and to have that mechanical energy be efficiently transferred from the piezoelectric medium to the MRR in the cladding. By shaping both media so that its joint mechanical resonance is at the microwave frequency, we can increase the efficiency of conversion from microwave to mechanical energy. In addition, we will want to choose a mechanical resonance with maximum oscillation in the vicinities of the piezoelectric medium and the part of the cladding containing the MRR.
@@
III.2 Theoretical model
Quantum-mechanically, we can describe the transduction (in the single-mode approximation) with the following Hamiltonian:
@@ III.2 Theoretical model
Not captured in the transduction Hamiltonian are the input, output, external microwave, and optical bosonic mode fields. As these fields are of bosons, we can use standard input-output relations [19, 20] and phenomenologically introduce loss and noise terms into the Heisenberg equations of motion describing the annihilation operators of these fields. These modified Heisenberg equations are known as Heisenberg-Langevin equations. In addition, we also assume that optical backscattering is negligible, so that we need not consider the reverse-propagating optical modes, which would double the complexity of the resulting dynamics for little benefit.
@@
Figure 3: Diagram showing the mode couplings in a quasi-schematic of a piezo-opto-mechanical transducer. The microwave mode $\hat{c}_{in}$ couples to the mechanical mode $\hat{b}$ piezoelectrically. The mechanical mode $\hat{b}$ couples to the optical mode $\hat{a}_{1}$ in the cladding MRR optomechanically. The external MRR optical mode $\hat{a}_{2}$ is coupled both to $\hat{a}_{1}$ and $\hat{a}_{out}$ through evanescent optical coupling interactions.
III.2.1 Consolidating the piezoelectric interaction
To obtain a simplified set of linear Heisenberg-Langevin equations of motion for the transducer (and from them, the transduction efficiency), we begin as was done in [16] by consolidating the equations of motion for the microwave and mechanical fields into one so that the mechanical field $\hat{b}$ can be regarded as being directly pumped by the microwave field $\hat{c}_{in}$ via the input-output relation (in frequency space):
$\tilde{c}_{in}+\tilde{c}_{out}=\sqrt{\gamma_{ex}}\tilde{b}$ (2)
Here, $\tilde{c}_{in}$ and $\tilde{c}_{out}$ are the annihilation operators in frequency space for photons traveling in the microwave transmission line, and $\gamma_{ex}$ is the external coupling constant between the microwave and mechanical fields. The overall coupling constant $\gamma_{ex}$ is given in terms of the microwave susceptibility $\chi_{MW}(\omega)$, the electromechanical (piezoelectric) coupling rate $g_{EM}$, and the coupling constant $\Gamma_{ex}$ between the microwave input and the microwave mode in the transducer:
$\sqrt{\gamma_{ex}}$ $=-ig_{EM}\chi_{MW}\sqrt{\Gamma_{ex}}$ (3)
$\frac{d\hat{c}}{dt}\!$ $=\!-i\omega_{MW}\hat{c}\!-\!\frac{\Gamma}{2}\hat{c}\!-\!ig_{EM}(\hat{b}\!+\!\hat{b}^{\dagger})\!+\!\!\sqrt{\Gamma_{0}}\hat{f}_{MW}\!+\!\!\sqrt{\Gamma_{ex}}\hat{c}_{in}$ (5b)
where $\hat{f}_{m}$ and $\hat{f}_{MW}$ are (bath) noise-mode annihilation operators for the mechanical and microwave fields, respectively.
In frequency space, the Heisenberg-Langevin equations become:
$\frac{1}{\chi_{MW}}\tilde{c}$ $=-ig_{EM}(\tilde{b}+\tilde{b}^{\dagger})+\sqrt{\Gamma_{0}}\tilde{f}_{MW}+\sqrt{\Gamma_{ex}}\tilde{c}_{in}$ (6b)
$\frac{d\hat{b}}{dt}=-i\omega_{m}\hat{b}-\frac{\gamma_{m}}{2}\hat{b}+\sqrt{\gamma_{ex}}\hat{c}_{in}+\sqrt{\gamma_{noise}}\hat{f}_{piezo}$ (7)
where:
$\frac{\gamma_{m}}{2}$ $=\frac{\gamma_{0}}{2}+g_{EM}^{2}\chi_{MW}$ (8a)
In addition to the input-output relation describing the consolidated piezoelectric interaction, we will also use an input-output relation to describe the coupling interaction of the pump field in the bus waveguide with the external MRR:
$\tilde{a}_{in}+\tilde{a}_{out}=\sqrt{\kappa_{ex,2}}\hat{a}_{2}$ (9)
Here, $\tilde{a}_{in}$ and $\tilde{a}_{out}$ are the annihilation operators (in frequency space) for the incoming and outgoing field modes in the bus waveguide, and $\kappa_{ex,2}$ is the optical coupling constant for light moving between the bus waveguide and its adjacent MRR.
To clarify future discussion we will assume the MRRs are identical (so that $\omega_{1}=\omega_{2}$). Moreover, we assume the pump laser is operating at frequency $\omega_{L}$, and transform to a frame of reference rotating with this frequency:
III.2.2 Linearizing the optomechanical interaction
Alternatively, if we set the pump laser frequency $\omega_{L}$ to be below $\omega_{1}$ by amount $\omega_{m}$ (so that $\Delta_{1}\approx\omega_{m}$), we end up in the transduction regime where the transduction portion of the interaction oscillates slowly, and the squeezing portion oscillates rapidly so that now it is the squeezing that does not contribute significantly to the system dynamics. Considering the system being driven in the two-mode squeezing regime to create entangled microwave-telecom photon pairs has its applications in quantum networking [22], but in what follows, we will consider the system exclusively in the transduction regime. In this regime, $\omega_{L}\approx\omega_{1}-\omega_{m}$, and we shall assume the bandwidth of the pump laser is narrow enough that we may neglect any two-mode squeezing contribution to the optomechanical interaction term of the system hamiltonian (12).
$\frac{1}{\chi_{02}(\omega)}\delta\tilde{a}_{2}$ $=iJ\delta\tilde{a}_{1}+\sqrt{\kappa_{0,2}}\tilde{f}_{0,2}+\sqrt{\kappa_{ex,2}}\delta\tilde{a}_{in}$ (13b)
$\frac{1}{\chi_{m}(\omega)}\tilde{b}$ $=i\bar{G}\bar{a}_{1}^{*}(\omega+\Delta_{t})\delta\tilde{a}_{1}+\sqrt{\gamma_{0}}\tilde{f}_{m}+\sqrt{\gamma_{ex}}\tilde{c}_{in}$ (13c)
$\tilde{a}_{in}+\tilde{a}_{out}=\sqrt{\kappa_{ex,2}}\tilde{a}_{2}$ (13d)
$\tilde{c}_{in}+\tilde{c}_{out}=\sqrt{\gamma_{ex}}\tilde{b}$ (13e)
$\frac{1}{\chi_{01}(\omega)}\tilde{\bar{a}}_{1}=iJ\tilde{\bar{a}}_{2}$ (14a)
$\frac{1}{\chi_{02}(\omega)}\tilde{\bar{a}}_{2}=iJ\tilde{\bar{a}}_{1}+\sqrt{\kappa_{ex,2}}\tilde{\bar{a}}_{in}$ (14b)
where we define the susceptibilities:
@@ III.2.2 Linearizing the optomechanical interaction
where the shift comes from incorporating the linear phase in the Fourier transform. Note that, from the context of the equations of motion, we have $\kappa_{1}=\kappa_{0,1}$; $\kappa_{2}=\kappa_{0,2}+\kappa_{ex,2}$; and $\gamma_{m}=\gamma_{0}+4g_{EM}^{2}/\Gamma$. To simplify things further, we will assume $\Delta_{t}\approx 0$ relative to the laser linewidth. With this, we have obtained the simplified equations of motion from which we will obtain the transduction efficiency.
III.3 Solving to find the transduction efficiency
With the equations of motion given as a linear system, we can solve them to obtain $\hat{a}_{out}$ as a linear function of $\hat{c}_{in}$, among other variables, and from this, the transduction efficiency.
$\delta\tilde{a}_{out}$ $\approx G_{\tilde{a}_{out},\tilde{c}_{in}}\tilde{c}_{in}+G_{\tilde{a}_{out},\tilde{a}_{in}}\delta\tilde{a}_{in}$
$+G_{\tilde{a}_{out},\tilde{f}_{m}}\tilde{f}_{m}+G_{\tilde{a}_{out},\tilde{f}_{0,1}}\tilde{f}_{0,1}$
$+G_{\tilde{a}_{out},\tilde{f}_{0,2}}\tilde{f}_{0,2}$ (16)
Here, the coefficient $G_{\tilde{a}_{out},\tilde{c}_{in}}$ is called the overall gain between modes $\tilde{c}_{in}$ and $\tilde{a}_{out}$, and is a transfer function in the context of Fourier analysis. As a function of $\omega$, its magnitude square is the transduction efficiency from microwave photons at frequency $\omega$, to optical photons at frequency $\omega+\omega_{L}$.
While one could compute $G_{\tilde{a}_{out},\tilde{c}_{in}}$ by brute force from the system of equations (III.2.2)(III.2.2), reference [16] introduces representing the system as a signal flow graph and using Mason’s Gain rule [23] in order to obtain the transduction efficiency in a simpler fashion that gives intuition on the internal dynamics of the system.
In Fig. 4, we show the signal flow graph for our system. With this signal flow graph, we can compute the transduction efficiency as the magnitude square of the overall gain from $\hat{c}_{in}$ to $\hat{a}_{out}$.
First, there is only one path from $\hat{c}_{in}$ to $\hat{a}_{out}$ with overall gain:
where the path goes from node $\hat{c}_{in}$ to $\hat{b}$ to $\hat{a}_{1}$ to $\hat{a}_{2}$ to $\hat{a}_{out}$ (following along the arrows), and the path gain is the product of the gains between each pair of adjacent nodes in the path. There are only two loops (which are defined as paths that begin and end on the same node without touching any other node twice), and they have gains:
$L_{1}$ $=-\bar{G}^{2}|\bar{a}_{1}|^{2}\chi_{01}\chi_{m}$ (18)
$L_{2}$ $=-J^{2}\chi_{01}\chi_{02}$ (19)
With this, Mason’s Gain formula gives for the overall gain from $\hat{c}_{in}$ to $\hat{a}_{out}$:
$\frac{\tilde{a}_{out}}{\tilde{c}_{in}}=\sum_{i}\frac{G_{i}\Delta_{i}}{\Delta}\;\;\rightarrow\;\;\frac{G_{1}\Delta_{1}}{\Delta}$ (21)
which in terms of our system variables is given by:
@@ III.3 Solving to find the transduction efficiency
$\boxed{\frac{\tilde{a}_{out}}{\tilde{c}_{in}}\equiv G_{\tilde{a}_{out},\tilde{c}_{in}}\!\!\!=\!\!\frac{\sqrt{\kappa_{ex,2}}\sqrt{\gamma_{ex}}\chi_{01}\chi_{02}\chi_{m}\bar{G}\bar{a}_{1}J}{1+\bar{G}^{2}|\bar{a}_{1}|^{2}\chi_{01}\chi_{m}+J^{2}\chi_{01}\chi_{02}}}$ (22)
@@
whose magnitude square is the microwave-optical transduction efficiency. Here we note that the amplitude for the inverse process of optical-microwave transduction is different due to using $\bar{a}_{1}^{*}$ instead of $\bar{a}_{1}$, but the magnitude square (i.e., the efficiency) is identical in both directions. While this may seem unusual, it is worth pointing out that the forward and reverse processes (i.e., direct vs converse piezoelectricity and photoelasticity vs electrostriction) are coupled with the same coefficients. This point will be elaborated on in the next section.
@@
Note: Upon close observation, one might conclude that if light coursing through either resonator should induce vibrations through the optomechanical interaction (i.e., electrostriction), then it is unrealistic to have an optomechanical interaction in the transducer Hamiltonian (III.2) acting on only the cladding micro-ring resonator. However, assuming that the acoustic modes associated to each MRR are decoupled from one another (e.g., are spatially separated), we may treat the optomechancial coupling in the external resonator as a loss mechanism (incorporated into $\kappa_{0,2}$) that may be incorporated into the Heisenberg-Langevin equations of motion. Moreover, by choice of design, we can minimize this ancillary optomechanical coupling by shaping the substrate around the second MRR so that it sits within a vibrational node at the appropriate mechanical frequency.
@@
In order to use this formula for the transduction efficiency amplitude, we will also need to know $\bar{a}_{1}(\omega)$, which can be solved as a function of the input pump spectrum $\bar{a}_{in}(\omega)$ using the equations (III.2.2):
$\boxed{\tilde{\bar{a}}_{1}=\frac{iJ\chi_{01}(\omega)\chi_{02}(\omega)\sqrt{\kappa_{ex,2}}}{1+J^{2}\chi_{01}(\omega)\chi_{02}(\omega)}\tilde{\bar{a}}_{in}}$ (23)
@@
Where the ratio $\tilde{\bar{a}}_{1}/\tilde{\bar{a}}_{in}$ is the amplitude gain of the field in the first MRR as a function of frequency (i.e., the amplitude gain spectrum), we will explore how the inter-MRR coupling $J$ causes a splitting in the amplitude gain spectrum in a later section. On resonance, this cavity enhancement factor simplifies to:
$\frac{|\tilde{\bar{a_{1}}}|^{2}}{|\tilde{\bar{a_{in}}}|^{2}}\approx\frac{64J^{2}\kappa_{ex,2}}{(\kappa_{1}+\kappa_{2})^{2}|(\kappa_{1}-\kappa_{2})^{2}-16J^{2}|}$ (24)
@@
For the system parameters used in [16], this cavity enhancement factor is actually much less than unity (about $2.6\times 10^{-9}$), though still sharply peaked compared to adjacent frequencies (with an effective $Q$ factor for this peak of $2.2\times 10^{6}$). However, this describes only the fraction of driving laser light on resonance that couples into $\hat{a}_{1}$, and does not compromise the overall transduction efficiency, where other inputs into $\hat{a}_{1}$ must be taken into account.
@@
IV Physical foundations of the interaction constants
IV.1 The Piezoelectric interaction
IV.1.1 Constitutive equations of piezoelectricity
The equations governing the piezoelectric interaction [24, 25] are determined from thermodynamics. The differential change in internal energy $dU$ of a dielectric (of which piezoelectric media are a subtype) can be given by the expression:
$dU=TdS+X_{ij}dx_{ij}+E_{i}dD_{i}$ (26)
$F$ $=U-TS$ (27a)
$dF$ $=-SdT+X_{ij}dx_{ij}+E_{i}dD_{i}$ (27b)

### v4

@@ Abstract
Two-way microwave-optical quantum transduction is essential to connecting distant superconducting qubits via optical fiber, and to enable quantum networking at a large scale. In Blésin, Tian, Bhave, and Kippenberg’s article, “Quantum coherent microwave-optical transduction using high overtone bulk acoustic resonances” (Phys. Rev. A, 104, 052601 (2021)), they lay out a two-way quantum transducer converting between microwave photons and telecom-band photons by way of an intermediary GHz-band phonon mode utilizing piezoelectric and optomechanical interactions respectively (and are the first to work out the quantum piezoelectric coupling). In this work, we examine both the piezoelectric, and optomechanical interactions from first principles, and together with the evanescent coupling between optical modes, discuss what parameters matter most in optimizing this kind of quantum transducer. For its additional utility, we have also compiled a table of relevant properties of optical materials that may be used as elements in transducers.
@@ I Introduction
Quantum transduction (i.e., where quantum information is transferred from one medium to another) is both a critical and fundamental process needed to connect different species of qubits in a quantum network. This makes more diverse heterogeneous quantum networks possible, and also allows us to leverage the complementary advantages [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] of different species of qubits. In particular, superconducting qubits, which function at microwave frequencies, enable rapid and deterministic gate operations, while photon-based qubits allow one to send quantum information long distances over optical fiber at ambient temperatures with relatively low loss (e.g., $0.18$dB/km for SMF28-Ultra 200 optical fiber at $1550$nm [11]) compared to sending microwave signals over coaxial cables (of the order $10^{2}$ to $10^{3}$dB/km). However, the (microwave/GHz scale) frequencies of light that interact best with superconducting qubits are five orders of magnitude below the (telecom) frequencies at which light propagates best in fiber optics. Moreover, the amount of random thermally generated microwave photons that happen to be in a propagating microwave mode compared to a single-photon microwave signal is large enough that resolving a transduced microwave signal from noise requires a cryogenic operating environment  [^1] As a figure of merit, the expected number of $3.285$ GHz microwave photons in a single mode at $298$K, assuming Boltzmann statistics, would be about $1.89\times 10^{3}$, compared to about $0.26$ at $100$mK.. Unlike classical networking, quantum signals cannot be simply amplified for long-distance transmission due to the no-cloning theorem [13]. Linear amplification of quantum signals (possibly written on individual photons) cannot be carried out without adding noise [14, 15]. This noisy amplification to compensate for loss eventually reaches an irreparable threshold where specific and distinct quantum states (e.g., non-Gaussian states) cannot be reliably transmitted over long distances. In order to connect distant superconducting qubits — as is needed to establish a network of superconducting quantum computers — it is necessary to coherently convert microwave-photons to telecom photons and back again with as little loss/noise as possible.
@@
In this work, we begin by discussing the essential elements underlying the piezoelectric and optomechanical interactions in order to determine what factors are most easily adjusted to maximize the efficiency of transduction. First, we give a brief overview of the system in question, showing how the transduction efficiency is obtained as a function of the coupling interactions. Following this, we review these coupling constants (derived ab initio in the appendices), and explore what material parameters and device design principles can be used to maximize both the transduction efficiency as well as the bandwidth over which efficient transduction can be accomplished, while minimizing noise.
II The Piezo-optomechanical transducer
II.1 The physical system
@@
II.2 Theoretical model of transduction efficiency
Quantum-mechanically, we can describe the transduction process (in the single-mode approximation) with the following Hamiltonian:
@@ II.2 Theoretical model of transduction efficiency
Not captured in the transduction Hamiltonian are the input, output, external microwave, and optical bosonic mode fields. As these fields are of bosons, we can use standard input-output relations [19, 20] to phenomenologically introduce loss and noise terms into the Heisenberg-Langevin equations for the fields within the transducer. In addition, we also assume that optical backscattering is negligible, so that we need not consider reverse-propagating optical modes, which would double the complexity of the resulting dynamics for little benefit.
For simplicity of discussion, we leave the detailed derivation of the transduction efficiency for Appendix B (also described in [16]). Starting from (II.2), we first consolidate the equations of motion for the microwave and acoustic fields into one that governs the piezoelectric interaction. Next, we linearize the optomechanical interaction on the assumption that the driving pump is bright enough that it is not noticeably depleted by transduction events. Following this, we transform to the frequency basis to convert the set of linear differential equations into linear algebraic equations, and use the mathematics of signal flow graphs to concisely obtain an amplitude for transduction.
@@
Figure 2: Diagram showing the mode couplings in a quasi-schematic of a piezo-opto-mechanical transducer. The microwave mode $\hat{c}_{\text{\text{in}}}$ couples to the mechanical mode $\hat{b}$ piezoelectrically. The mechanical mode $\hat{b}$ couples to the optical mode $\hat{a}_{1}$ in the cladding MRR optomechanically. The external MRR optical mode $\hat{a}_{2}$ is coupled both to $\hat{a}_{1}$ and $\hat{a}_{\text{out}}$ through evanescent optical coupling interactions.
Defining the susceptibilities:
@@
we can express the transduction amplitude as:
@@
$\boxed{\frac{\tilde{a}_{\text{out}}}{\tilde{c}_{\text{in}}}\equiv G_{\tilde{a}_{\text{out}},\tilde{c}_{\text{in}}}\!\!\!=\!\!\frac{\sqrt{\kappa_{ex,2}}\sqrt{\gamma_{ex}}\chi_{01}\chi_{02}\chi_{m}\bar{G}\bar{a}_{1}J}{1+\bar{G}^{2}|\bar{a}_{1}|^{2}\chi_{01}\chi_{m}+J^{2}\chi_{01}\chi_{02}}}$ (3)
where $\gamma_{m}$ is the overall mechanical linewidth of the bulk acoustic resonator, $\kappa_{1}$ is the resonance linewidth of the first MRR, and $\kappa_{2}$ is the resonance linewidth of the second MRR (which is generally different from $\kappa_{1}$ due to the presence of the coupling to the bus waveguide).
@@
The magnitude square of $G_{\tilde{a}_{\text{out}},\tilde{c}_{\text{in}}}$ is the microwave-optical transduction efficiency. Here we note that the amplitude for the inverse process of optical-microwave transduction is different due to using $\bar{a}_{1}^{*}$ instead of $\bar{a}_{1}$, but the magnitude square (i.e., the efficiency) is identical in both directions. While this may seem unusual, it is worth pointing out that the forward and reverse processes (i.e., direct vs converse piezoelectricity and photoelasticity vs electrostriction) are coupled with the same coefficients (See Appendix C for details).
@@
Note: Upon close observation, one might conclude that if light coursing through either MRR should induce vibrations through the optomechanical interaction (i.e., electrostriction), then it is unrealistic to have an optomechanical interaction in the transducer Hamiltonian (II.2) acting on only the cladding micro-ring resonator. However, assuming that the acoustic modes associated to each MRR are decoupled from one another (e.g., are spatially separated), we may treat the optomechancial coupling in the external resonator as a loss mechanism (incorporated into $\kappa_{0,2}$) that may be incorporated into the Heisenberg-Langevin equations of motion. Moreover, by choice of design, we can minimize this ancillary optomechanical coupling by shaping the substrate around the second MRR so that it sits within a vibrational node at the appropriate mechanical frequency.
@@
In order to use this formula for the transduction efficiency amplitude, we will also need to know $\bar{a}_{1}(\omega)$, which can be solved as a function of the input pump spectrum $\bar{a}_{\text{in}}(\omega)$ using the equations (52):
$\boxed{\tilde{\bar{a}}_{1}=\frac{iJ\chi_{01}(\omega)\chi_{02}(\omega)\sqrt{\kappa_{ex,2}}}{1+J^{2}\chi_{01}(\omega)\chi_{02}(\omega)}\tilde{\bar{a}}_{\text{in}}}$ (4)
@@
Where the ratio $\tilde{\bar{a}}_{1}/\tilde{\bar{a}}_{\text{in}}$ is the amplitude gain of the field in the first MRR as a function of frequency (i.e., the amplitude gain spectrum), we will explore how the inter-MRR coupling $J$ causes a splitting in the amplitude gain spectrum in a later section. On resonance, this cavity enhancement factor simplifies to:
$\frac{|\tilde{\bar{a_{1}}}|^{2}}{|\tilde{\bar{a_{\text{in}}}}|^{2}}\approx\frac{64J^{2}\kappa_{ex,2}}{(\kappa_{1}+\kappa_{2})^{2}|(\kappa_{1}-\kappa_{2})^{2}-16J^{2}|}$ (5)
@@
III Physical foundations of the interaction constants
III.1 The Piezoelectric interaction
III.1.1 Constitutive equations of piezoelectricity
The equations governing the piezoelectric interaction [21, 22] are determined from thermodynamics. For a full discussion of their derivation, see Appendix C.
Using standard definitions for the inverse dielectric permittivity tensor $\eta_{ij}$, the elasticity or stiffness tensor $c_{ijk\ell}$, and the stress-voltage form of the piezoelectric tensor $h_{ijk}$  [^3] Note that the more common stress-charge form of the piezoelectric tensor is $e_{ijk}$ such that $h_{ijk}=\eta_{im}e_{mjk}$., and having assumed a negligible difference between the isothermal and isentropic forms of these constants [24], we obtain the constitutive equations for the piezoelectric interaction: